In an isosceles triangle ABC, Angle C is the vertex angle. If angle A is 3x-6 and angle C is 2x, find the measure of angle B.
Thank you for helping me!

Answer by stanbon(75887)

In an isosceles triangle ABC, Angle C is the vertex angle. If angle A is 3x-6 and angle C is 2x, find the measure of angle B.
-------
A + B + C = 180 deg and A = B
------
3x-6 + 3x-6 + 2x = 180
8x = 192
x = 24
----
B = 3x-6 = 3*24-6 = 66 degrees
-----------
Cheers,
Stan H.
-----------